# Dynamic Neural Potential Field: Online Trajectory Optimization in the Presence of Moving Obstacles

### v2

@@ 3 Background
Explicitly predicting obstacle trajectories can provide precise future states but introduces (i) an additional module to train and maintain, (ii) error compounding over the horizon, and (iii) coupling between prediction errors and MPC feasibility. Our approach estimates repulsive potentials that directly shape the MPC cost, offering two key benefits: (a) the optimization remains robust to small prediction inaccuracies because the potential encodes spatial risk rather than a single hypothesized pose, and (b) potentials are differentiable and integrate naturally with L4CasADi for efficient gradient computation. When high-quality motion predictors are available, D1 can consume predicted trajectories in the form of future obstacle maps, while D2/NPField-GPT can be conditioned on predictor states $I_{dyn}=(x,y,\theta)$ and use only the current obstacle map, implicitly extracting information about the dynamic obstacle’s footprint. Thus, our framework complements—rather than replaces—modern predictors by converting them into optimization-friendly costs.
4 Neural Network-Based Potential Field Generation
We study three variants—NPField-D1, D2, and NPField-GPT — for producing footprint-aware repulsive potentials in dynamic scenes. In all cases, a lightweight map/footprint encoder runs once per replan to produce a spatial embedding; at query time, the robot pose and dynamic-obstacle state condition the potential predictor. Full architectural diagrams are shown in Fig. 2. Detailed descriptions of D1/D2 and the auxiliary reconstruction loss are deferred to Appendix C.
Why GPT for dynamic potentials?
NPField-GPT replaces independent per-step heads with an autoregressive Transformer that predicts a short horizon of potentials as a sequence. Causal self-attention conditions each step t+1t{+}1 on all prior steps, improving temporal consistency and safety margins relative to one-shot predictions. In our setting, this offers: (i) temporal coherence under motion changes and occlusions; (ii) long-range cross-step conditioning between map features and obstacle motion; (iii) compatibility with MPC via exact differentiation (L4CasADi) and a compact token interface; and (iv) low amortized latency, since the map/footprint encode is computed once and reused.
Concretely, NPField-GPT tokenizes four inputs—map/footprint embedding, robot-conditioned map features, dynamic-obstacle state, and the query pose—and autoregressively emits a length-10 sequence of scalar potentials, one per horizon step. Each new prediction is fed back as a token, enabling consistent, risk-aware forecasts in cluttered corridors.
Baselines.
D1 treats a dynamic scene as a sequence of static frames evaluated step-wise; D2 uses parallel MLP heads to predict potentials at all steps from the initial state. These provide strong speed/parameter baselines but lack NPField-GPT’s temporal coupling; see Appendix C.
Figure 2: NPField-D1 (left), NPField-D2 (middle), and NPField-GPT (right). Yellow: encoder (map and footprint); green: per-sample potential predictor; red: auxiliary decoder used only for training. D2 uses parallel heads for $t_{0},\dots,t_{10}$, whereas NPField-GPT is an autoregressive transformer-based model that conditions t+1t{+}1 on all prior outputs.
5 Dataset preparation
We construct a supervised dataset of footprint-aware repulsive potentials on small occupancy-grid crops. Each example consists of: (i) a 50×50 occupancy sub-map at 0.1 m/cell, cropped from MovingAI city maps [38] and an in-house office grid; (ii) the Husky UGV footprint (two variants: folded arm and outstretched arm); (iii) a dynamic obstacle state (0.7×0.5 m rectangle with pose $(x,y,\theta)$ and constant-velocity heading); and (iv) a query robot pose. The target is the footprint-aware repulsive potential computed from the signed distance field as $J_{o}=w_{1}(\pi/2+\arctan(w_{2}-w_{2}\,\mathrm{SDF}))$, evaluated as the maximum over cells covered by the footprint. Full generation protocol, dataset counts, and splits are provided in Appendix A.
6 Experiments
Implementation summary. We use Acados (SQP) [39] with CasADi [40] and integrate NPField via L4CasADi [33] for exact differentiation of the neural cost; the stack runs on an AMD Ryzen 5 3500 CPU and an NVIDIA GeForce RTX 2080. End-to-end replanning is typically 0.35–0.70 s across D1,D2,NPField-GPT, dominated by the MPC solve; we pair the local planner with a Theta* [3] global planner. Full implementation, runtime, and scalability details are provided in Appendix B.
Figure 3: Trajectory predictions for dynamic obstacles generated by NPField-D2 (middle row) and NPField-GPT (bottom row) under three distinct scenarios. Each column corresponds to a scenario, with the first, eighth, and tenth time steps visualized. The models predict repulsive potentials using only the initial map configuration (top row), while subsequent steps are autoregressively generated by NPField-GPT or predicted in parallel by D2 (via MLP heads). The rightmost scenario tests robustness by reversing the dynamic obstacle’s direction ($I_{dyn}$) relative to the validation dataset. Visualizations depict potentials across the full map for clarity, though the models inherently operate on per-query point-wise predictions.
Figure 4: Illustration of a dynamic obstacle navigation scenario using the NPField-GPT architecture. The left panels depict the initial environment (left) and the corresponding neural potential field (right), while the right panels show the final state after trajectory execution. NPField-GPT autonomously predicts obstacle motion based solely on the initial map, without requiring explicit obstacle trajectory inputs, demonstrating self-contained dynamic collision avoidance capabilities.
6.1 Numerical experiments
@@
We evaluated our algorithm on 100 scenarios (Fig. 4) using the [41] framework, which includes tasks like navigating through narrow passages. The assessment covered standard metrics such as planning time, path length, smoothness, and angle-over-length, where a lower value is preferable for all. Additionally, we introduced a custom metric, ”safety distance” (the minimum value of the SDF).
The experimental results, as detailed in table 1, compared three versions of the Dynamic NPField algorithm: D1, D2, and NPField-GPT, as well as MPPI and the CIAO* [10, 11] trajectory optimization algorithm. As demonstrated in Figure 3, both NPField-D2 and NPField-GPT effectively predict spatiotemporal repulsive potentials for dynamic obstacles. While D2 generates parallel predictions for multiple time steps using MLP heads, GPT employs an autoregressive GPT architecture to sequentially refine potentials, as shown in the eighth and tenth time steps. The rightmost scenario highlights the models’ ability to adapt to abrupt changes in obstacle motion, validating the fidelity of their dynamic embeddings ($I_{dyn}$).
@@
For each waypoint of the reference CIAO* approximate free space around the robot with a convex figure (circle or rectangle) and constrain the robot to be inside this figure. NPField-D2 and NPField-GPT directly use the data about the bounding box and motion direction of the dynamic obstacle, while NPField-D1 and CIAO* require the projection of the predicted bounding boxes onto the obstacle map.
@@
Overall, CIAO* showed average performance, which aligns with its status as a SOTA method. MPPI had similar results, where it was faster than CIAO* but failed to find optimal trajectories multiple times and overall had sub-par results. NPField-D1 was the fastest, registering a computation time of 359.51 ms, which is nearly 300 ms faster than CIAO*. In contrast, NPField-GPT opted for the shortest path and safest trajectory, as shown by the highest safety distance and lowest AOL, though at the cost of increased computation time and path length. This behavior, although anticipated due to the use of the GPT model, suggests that NPField-GPT might be the most appropriate choice for real-world scenarios where some trade-off in time and length is acceptable to enhance robot safety.
Table 1: Comparative studies on Dynamic obstacles scenarios
@@
NPField-D1 366.62 2.76 0.013 0.033 0.093
NPField-D2 493.50 2.15 0.017 0.058 0.109
NPField-GPT 663.81 2.06 0.011 0.015 0.115
6.2 Real robot experiments
@@
We tested the concept using a Husky UGV mobile manipulator (see Figure 5) as a ROS module. For local planning and control, we utilized MPC, and for global planning, we used the Theta* planner [3]. Our testing scenario involved the robot maneuvering through a complex map. The hardware specifications included an AMD Ryzen 5 3500 CPU and an NVIDIA GeForce RTX 2080 with CUDA support for PyTorch and L4CASADI. The entire navigation stack included a Cartographer and RTAB-Map for the global planner, with a second RTAB-Map for the local planner. All components of the control system were implemented as ROS nodes, with a central node managing communication with the Husky UGV hardware. Real-time planning and execution can be viewed in the supplementary video.
Figure 5: Husky UGV mobile manipulator
7 Conclusion, limitations and future work
@@
This article presents a novel approach to MPC collision avoidance, which extends Neural Potential Field for the case of an environment with dynamic obstacles. We propose three neural architectures (NPField-D1, NPField-D2, NPField-GPT) for predicting dynamic neural potential. We implement them and compare them with each other and with MPPI and CIAO* trajectory optimizers. NPField-D1 (modeling dynamic environment with a reference to static maps) and NPField-D2 (predicting a reference of potentials at once) outperform CIAO* and MPPI in path length and computation time. NPField-GPT (autoregressive potential prediction) is slower; however, it shows better potential reconstruction in general.
@@
In common, performance measures from table 1 show that our controllers require hundreds of milliseconds to replan the trajectory. This is sufficient for an indoor mobile robot with limited velocity and 1 Hz replanning rate, while faster systems such as cars or drones require a higher rate. Performance may be improved by using more powerful hardware and developing faster models and implementations. Numerical MPC, in general, is sensitive to the quality of the initial guess. In general, it is recommended to use it as a local planner together with a global planner, which provides a rough but suboptimal reference path. Also, it is worth noting that NPField is a learning-based method and therefore it has no theoretical safety guarantees. Therefore, the obtained solution has to be checked for collisions before its execution (which may be easily done by projecting the footprint onto the optimized waypoints).
@@
A significant assumption for the MPC problem is that for each dynamic obstacle, we know either a prediction of its future trajectory (NPField-D1) or a constant direction of its movement. We compensate for the inaccuracy of this assumption via replanning the trajectory on each step. A meaningful direction of future research consists of overcoming this limitation by integrating autoregressive potential forecasting with advanced motion prediction techniques.
Appendix A Dataset generation details
Maps and crops.
We crop large occupancy grids into 50 $\times$ 50 sub-maps at 0.1 m/cell (5 $\times$ 5 m windows), yielding 1000 unique sub-maps for data generation (Fig. 6).
Figure 6: Cropping original maps into sub-maps
Dynamic obstacle placements.
For each sub-map, we sample one agent position with three orientations. Assuming 0.3 m/s constant speed, we generate 10 forward positions per orientation (31 variants per sub-map: 1 static-only; 3 $\times$ 10 with the agent).
Reference potentials.
For each sub-map copy, we compute the signed distance field (SDF) and derive cell-wise potentials as $J_{o}=w_{1}(\pi/2+\arctan(w_{2}-w_{2}\,\mathrm{SDF}))$. For two Husky UGV footprints (folded and outstretched arm), we evaluate 10 random orientations of the footprint at each query pose and take the maximum over covered cells as the footprint-aware target.
Dataset structure and splits.
The dataset is tabular with fields: sub-map ID, query robot pose, dynamic obstacle pose/direction, footprint id, and the target potential. We split by sub-map into train/val/test as 70/15/15%. Validation figures (e.g., Fig. 3) use held-out sub-maps and placements. We also include stress tests by reversing $I_{dyn}$ on validation sub-maps.
Appendix B Implementation details, runtime, and scalability
Solver and toolchain.
We solve nonlinear MPC with SQP (Acados [39]) and CasADi [40]. We integrate PyTorch NPField models using L4CasADi [33] for exact symbolic differentiation (preferred over ML-CasADi [42], which uses local Taylor approximations).
Neural integration.
The map/footprint encoder runs once per replan on GPU; point-wise potential heads (D1/D2) are batched over horizon steps, while NPField-GPT runs autoregressively across time steps.
Runtime breakdown.
On AMD Ryzen 5 3500 CPU and NVIDIA RTX 2080: (i) map/footprint preprocessing and $h_{\text{map}}$ (GPU) 6–12 ms; (ii) per-sample embedding and NPField forward (GPU) D1: 30–40 ms, D2: 55–80 ms, NPField-GPT: 120–160 ms; (iii) Acados SQP solve with L4CasADi cost (CPU) 250–550 ms; (iv) ROS I/O 10–20 ms.
Scalability.
Let $N$ be horizon and $M$ dynamic obstacles. D1: $\mathcal{O}(N)$ single-step calls; D2: $\mathcal{O}(1)$ head depth per step (batched); NPField-GPT: $\mathcal{O}(N)$ autoregression with shared attention. MPC is roughly quadratic in $N$ with dense costs; we use warm starts and structure. For multiple obstacles, we concatenate $I_{dyn}$ and cap $M$ by the nearest obstacles per window.
Planning stack.
We pair the local MPC with Theta* [3] for global polylines. Theta* uses a circular footprint for robustness; local MPC enforces footprint-aware safety.
Appendix C Architectural details and auxiliary losses
NPField-D1 (static-frame decomposition).
A dynamic scene is represented as a sequence of static maps; the potential is evaluated per step with the same spatial encoder. This yields strong latency and simplicity, but temporal consistency is implicit and delegated to the MPC horizon.
NPField-D2 (parallel heads).
A single encoder is paired with multiple lightweight MLP heads, one per future step $(t_{0},\dots,t_{10})$, conditioned on the dynamic-obstacle state $I_{dyn}=(x,y,\theta)$. This keeps parameter count low and supports batched prediction, but steps are independent and may drift temporally under distribution shift.
Auxiliary reconstruction loss.
During training only, we decode the map embedding back to an occupancy grid with a small CNN head and apply binary cross-entropy over cells. This regularizes the spatial embedding, improving stability for NPField-D2/NPField-GPT without affecting test-time runtime (the decoder is discarded).
Appendix D Training hyperparameters and pseudocode
Model hyperparameters (NPField-GPT head).
  • Transformer: $d_{model}{=}864$, $n_{head}{=}8$, $n_{layer}{=}6$, dropout $0.1$; autoregressive causal mask; MLP output head.
  • Map/footprint encoder: hidden channels $64$, attention blocks $4$, attention heads $4$, CNN dropout $0.15$, downsample steps $2$, input $50\times 50$.
  • Decoder projection: flattened feature size $3456$ $\rightarrow$ $d_{model}$.
  • Outputs: horizon length $10$; loss is MSE on normalized potentials.
Optimization.
  • Optimizer: Adam, learning rate $1\times 10^{-4}$, gradient clipping at $0.9$.
Dataset interface.
Each sample contains: (i) stacked map and footprint, (ii) query pose $(x,y,\theta)$, (iii) dynamic obstacle state $(x,y,\theta)$, and (iv) a length-10 potential target sequence for the scenario window.
Training loop (pseudocode).
  1. 1.
    Build dataloaders with the fixed dataset covering all map/pose/head combinations.
  2. 2.
    For each batch: encode map/footprint once; encode query pose and dynamic state; run autoregressive Transformer to predict a length-10 potential sequence.
  3. 3.
    Compute MSE to the target sequence; backprop, clip gradients, step optimizer.
Appendix E MPC integration pseudocode
  1. 1.
    Wrap the trained PyTorch potential head with L4CasADi to obtain a differentiable CasADi function that maps $\text{embedding},x,y,\theta]\mapsto$ horizon potentials.
  2. 2.
    In the Acados model, define states $(x,y,v,\theta,t)$ and controls $(a,\omega,T)$; pass the embedding as a parameter vector.
  3. 3.
    Inside the cost, call the L4CasADi function, flatten the returned vector, and select the element matching the current time slice.
  4. 4.
    Minimize a nonlinear least-squares objective combining tracking, control effort, and the neural potential term; solve with SQP.

### v3

@@ III Background
Explicitly predicting obstacle trajectories can provide precise future states but introduces (i) an additional module to train and maintain, (ii) error compounding over the horizon, and (iii) coupling between prediction errors and MPC feasibility. Our approach estimates repulsive potentials that directly shape the MPC cost. This has two benefits: (a) optimization remains robust to small prediction errors because potential encodes spatial risk rather than a single hypothesized obstacle pose, and (b) potentials are differentiable and integrate naturally with L4CasADi for efficient gradient computation. When high-quality motion predictors are available, StaticMLP can consume predicted trajectories in the form of future obstacle maps, while DynamicMLP/NPField-GPT can be conditioned on predictor states $I_{dyn}=(x,y,\theta)$ and infer horizon risk directly from the current map context. Thus, our framework complements modern predictors by converting their outputs into optimization-friendly costs.
IV Neural Network-Based Potential Field Generation
We study three variants for producing footprint-aware repulsive potentials in dynamic scenes: NPField-StaticMLP, NPField-DynamicMLP, and NPField-GPT. All variants share a modular pipeline with three parts: (i) map/footprint encoding (yellow block), (ii) point-wise potential prediction for MPC query states (green block), and (iii) an auxiliary map-reconstruction decoder used only in training (red block). The auxiliary decoder regularizes the spatial embedding and is removed at inference time.
NPField-StaticMLP models dynamic navigation as a sequence of static maps and evaluates one-step potentials for each frame. This keeps the predictor simple but requires explicit future-map construction from obstacle forecasts.
NPField-DynamicMLP predicts all horizon steps in parallel with separate MLP heads conditioned on dynamic obstacle state $I_{dyn}=(x,y,\theta)$. This is efficient but temporal coupling between steps is weak because each step-specific head is independent.
NPField-GPT uses a Transformer backbone with non-autoregressive horizon prediction. Instead of generating step t+1t{+}1 from step $t$, NPField-GPT appends learnable query tokens (10 for the horizon) and predicts the full potential sequence in a single forward pass.
Token sequence.
The input to the Transformer is a single sequence of tokens. The context is built from 11 tokens in a fixed order: (1) map embedding (CNN plus spatial self-attention over the occupancy grid, projected to dimension $d$); (2) robot footprint embedding; (3)–(4) dynamic-obstacle position $(x_{dyn},y_{dyn})$ each encoded by a linear layer; (5)–(6) dynamic-obstacle orientation $\theta_{dyn}$ as $\sin\theta_{dyn}$ and $\cos\theta_{dyn}$; (7)–(8) query robot position $(x,y)$; (9)–(10) query robot orientation $\theta$ as $\sin\theta$ and $\cos\theta$; (11) a learned relative-coordinate fusion token that combines robot and obstacle coordinates via an MLP (context plus delta terms) and trainable gains to emphasize obstacle-relative geometry. The map and footprint encodings are computed once from the occupancy sub-map and footprint image; the coordinate tokens depend on the query pose $(x,y,\theta)$ and dynamic state $I_{dyn}=(x_{dyn},y_{dyn},\theta_{dyn})$. After the context, $H=10$ learnable query vectors are appended (the vectors themselves are fixed parameters, not a function of $(x,y,\theta)$). Learned positional embeddings are added to the full sequence; the stack is passed through the causal Transformer and layer norm. Because each query token attends over all preceding context tokens—including the robot pose $(x,y,\theta)$ and dynamic-obstacle state—the output at the query positions is conditioned on the query pose. The hidden states at the query positions are projected by a linear head and passed through sigmoid to obtain the predicted potential at each horizon step. Thus the potential horizon is predicted in one forward pass, with dependence on the query pose arising through attention to the context.
Training loss.
Let $\hat{y}_{b,h}$ and $y_{b,h}$ denote the predicted and target potential for batch index $b$ and horizon step $h$. Let $d_{b}$ be the Euclidean distance from the query pose $(x_{b},y_{b})$ to the dynamic obstacle $(x_{dyn,b},y_{dyn,b})$. We use a distance-dependent weight $w(d)=1+\alpha\exp\bigl(-(d/\sigma)^{2}/2\bigr)$ with $\alpha=2$, $\sigma=0.5$ to upweight errors near the obstacle. The potential loss is the weighted mean squared error
$\mathcal{L}_{\mathrm{pot}}=\frac{1}{BH}\sum_{b,h}w(d_{b})\,(\hat{y}_{b,h}-y_{b,h})^{2}.$ (2)
An auxiliary occupancy reconstruction loss $\mathcal{L}_{\mathrm{map}}$ is computed by decoding the hidden state at the last context token (the relative-coordinate fusion token) through a small decoder and applying cross-entropy to the ground-truth occupancy grid; this stabilizes the spatial embedding. The total training loss is
$\mathcal{L}=\mathcal{L}_{\mathrm{pot}}+\lambda_{\mathrm{map}}\,\mathcal{L}_{\mathrm{map}},$ (3)
with $\lambda_{\mathrm{map}}\geq 0$ (default 1). The auxiliary decoder and $\mathcal{L}_{\mathrm{map}}$ are used only during training and removed at inference.
V Dataset preparation and generation details
[…]
Figure 3: Cropping original maps into sub-maps
  1. 1.
    The maps are cropped into sub-maps and represented as images with $\times50$ resolution (Fig. 3), where each pixel corresponds to 0.1 m ($\times10$ cm) in the real environment (i.e., each sub-map covers $\times5$ m). This yields 1000 unique sub-maps for data generation.
  2. 2.
[…]
  3. 3.
    Every sub-map copy is transformed into a costmap where each cell is filled with the reference potential.
    1. (a)
      The signed distance function (SDF) is computed for each cell (distance to the nearest obstacle border; positive in free space, negative in obstacles).
    2. (b)
      Repulsive potential per cell is $J_{o}=w_{1}(\pi/2+\arctan(w_{2}-w_{2}\,\mathrm{SDF}))$. This sigmoid is low far from obstacles, approaches $w_{1}$ asymptotically inside obstacles, and has maximum derivative on the obstacle border.
  4. 4.
    Two robot footprints are used, both corresponding to a real Husky UGV mobile manipulator (folded and outstretched arm).
  5. 5.
    For each footprint, the potential at each query pose is obtained by placing the footprint at 10 random orientations per pixel for each copy of every sub-map; the robot footprint is projected onto the map and the maximum potential over the covered cells is taken as the repulsive potential.
The resulting dataset is tabular with fields: sub-map ID, query robot pose, dynamic obstacle pose/direction, footprint ID, and target potential.
Data splits and validation. We split the dataset into training/validation/test sets by sub-map (70/15/15%). All figures that reference a validation configuration use sub-maps and agent placements not present in training. We also include direction-reversal stress tests by flipping $I_{dyn}$ on held-out validation sub-maps to evaluate robustness.
VI Implementation and computational profile
We solve the nonlinear MPC problem via Sequential Quadratic Programming (SQP) using Acados [38] with CasADi [4] for differentiation. To integrate our PyTorch [28] potential model into the Acados solver, we wrap it using L4CasADi [31], which provides an exact symbolic description suitable for accurate gradients (unlike local approximations such as ML-CasADi [32]). This allows us to directly evaluate the predicted horizon potential in the Acados model, optimizing a nonlinear least-squares objective that combines tracking, control effort, and the neural potential term.
End-to-end runtime breakdown.
On an AMD Ryzen 5 3500 CPU and NVIDIA GeForce RTX 2080 (CUDA), typical per-replan times are:
  • map/footprint preprocessing and $h_{\text{map}}$ encoding (GPU): 6–12 ms;
  • per-sample point embedding and NPField forward over the horizon (GPU): NPField-StaticMLP: 30–40 ms, NPField-DynamicMLP: 55–80 ms, NPField-GPT: 120–160 ms (Transformer with parallel query-token horizon prediction);
  • Acados SQP solve with L4CasADi cost (CPU): 250–550 ms depending on variant and scenario complexity;
  • ROS I/O and sensor updates: 10–20 ms.
This yields end-to-end medians consistent with Table I. The bottleneck remains the MPC solve, while NPField-GPT adds moderate overhead due to the larger Transformer backbone and enhanced coordinate conditioning.
Scalability.
Let $N$ be the horizon length and $M$ the number of dynamic obstacles. For NPField-StaticMLP, complexity scales as $\mathcal{O}(N)$ forward calls of a single-step potential; for NPField-DynamicMLP, a constant number of heads gives $\mathcal{O}(1)$ forward depth per time step (batched over $N$); for NPField-GPT, horizon prediction is parallel in one Transformer pass, with complexity dominated by attention over context and query tokens. MPC cost is dominated by the nonlinear solve and scales roughly quadratically in $N$; we use warm starts and structured costs. With multiple obstacles, $I_{dyn}$ and token counts grow linearly in $M$; we cap $M$ per local window and prioritize nearest obstacles.
Training.
For NPField-GPT we use typical Transformer settings: 4 layers, 4 heads, embedding size 576; 10 learnable horizon query vectors; map encoder output projected to the model embedding width before Transformer fusion. Optimization uses an Adam-family optimizer with gradient clipping and optional mixed precision.
Our local planner works together with Theta* [25] global planner, which generates global plans as polylines. Note that Theta* uses a simplified version of the robot footprint (a circle with a diameter equal to the robot width) as it fails to provide a safe path with a complete footprint model. This simplified model does not guarantee the safety of the global plan; therefore, the safety of the trajectory is provided by our local planner.
VII Experiments
Figure 4: First and last frames of a representative validation scenario for each method. The heatmap encodes the predicted repulsive potential. The larger rectangle denotes the robot footprint; the smaller rectangle denotes the dynamic obstacle.
Figure 5: Five sequential time steps of MPC trajectory optimization. The sequence illustrates how NPField-GPT adapts the full-horizon potential in real time, enabling smooth collision avoidance without explicit future-map rollout.
VII-A Numerical experiments
@@
We evaluated our algorithm on 100 scenarios using the [10] framework, which includes tasks like navigating through narrow passages. Figure 5 shows a representative example: five successive MPC replan steps in which NPField-GPT steers the robot around a moving obstacle by continuously updating the predicted potential horizon. The assessment covered standard metrics such as planning time, path length, smoothness, and angle-over-length, where a lower value is preferable for all. Additionally, we introduced a custom metric, “safety distance” (the minimum value of the SDF).
The experimental results, as detailed in Table I, compared three versions of the Dynamic NPField algorithm: StaticMLP, DynamicMLP, and NPField-GPT, as well as MPPI and the CIAO* [33, 34] trajectory optimization algorithm. Figure 4 compares the predicted potential fields at the first and last frames of a validation scenario. The qualitative differences are clear: StaticMLP tends to merge the dynamic obstacle into the surrounding wall, losing its distinct boundary because the obstacle footprint is small relative to the map. DynamicMLP produces overly diffuse boundaries and does not always capture the temporal displacement of the obstacle correctly. NPField-GPT yields the sharpest and most accurate potentials, maintaining well-defined boundaries for both static walls and the moving obstacle throughout the horizon.
@@
For each waypoint of the reference, CIAO* approximates the free space around the robot with a convex figure (circle or rectangle) and constrains the robot to be inside this figure. DynamicMLP and NPField-GPT directly use the data about the bounding box and motion direction of the dynamic obstacle, while StaticMLP and CIAO* require the projection of the predicted bounding boxes onto the obstacle map.
@@
Overall, CIAO* showed average performance, which aligns with its status as a state-of-the-art method. MPPI was faster than CIAO* but frequently failed to find optimal trajectories, resulting in generally inferior performance. StaticMLP was the fastest, registering a computation time of 366.62 ms, which is over 400 ms faster than CIAO*. In contrast, NPField-GPT achieved the safest trajectory with the shortest path, as shown by the highest safety distance, lowest path length, and lowest AOL, though at the cost of increased computation time. This behavior suggests that NPField-GPT is the most appropriate choice for real-world scenarios where computation time can be traded for enhanced safety and path efficiency. In Table I, we abbreviate planner names as NPF-SMLP, NPF-DMLP, and NPF-GPT.
TABLE I: Comparative studies on Dynamic obstacles scenarios
@@
NPF-SMLP 366.62 2.76 0.013 0.033 0.093
NPF-DMLP 493.50 2.15 0.017 0.058 0.109
NPF-GPT 663.81 2.06 0.011 0.015 0.115
Figure 6: Example of multi-obstacle scenarios solved by NPField-GPT. The potential field is obtained by summing per-obstacle predictions from the single-obstacle model without retraining.
VII-B Extension to multiple dynamic obstacles
NPField-GPT naturally extends to scenes with multiple dynamic obstacles via additive superposition of repulsive potentials, requiring no architectural changes or retraining. Given $M$ dynamic obstacles, the combined occupancy map (with all obstacles rendered at the current timestep) is encoded once, while the per-obstacle dynamic state $I_{dyn}^{(j)}=(x_{dyn}^{(j)},y_{dyn}^{(j)},\theta_{dyn}^{(j)})$ is supplied individually to produce $M$ independent embeddings. At each MPC query point, the model is evaluated $M$ times and the resulting horizon potentials are summed: $J_{o}^{\text{multi}}=\sum_{j=1}^{M}J_{o}^{(j)}$. This superposition is integrated into the Acados solver via a single L4CasADi wrapper that loops over embeddings internally. We evaluate this scheme on scenarios with 2–5 moving obstacles (Fig. 6).
VII-C Real robot experiments
@@
We tested the concept using a Husky UGV mobile manipulator as a ROS module. For local planning and control, we utilized MPC, and for global planning, we used the Theta* planner [25]. Our testing scenario involved the robot maneuvering through a complex map. The entire navigation stack included a Cartographer [11] and RTAB-Map [19] for the global planner, with a second RTAB-Map for the local planner. All components of the control system were implemented as ROS nodes, with a central node managing communication with the Husky UGV hardware. Real-time planning and execution can be viewed in the supplementary video.
VIII Conclusion, limitations and future work
@@
This article presents a novel approach to MPC collision avoidance, which extends Neural Potential Field for the case of an environment with dynamic obstacles. We propose three neural architectures (NPField-StaticMLP, NPField-DynamicMLP, NPField-GPT) for predicting dynamic neural potentials and integrate them with real-time MPC.
@@
Overall, the performance measures in Table I show that our controllers require hundreds of milliseconds to replan the trajectory. This is sufficient for an indoor mobile robot with limited velocity and a 1–2 Hz replanning rate, while faster systems such as cars or drones require a higher rate. Performance may be improved by using more powerful hardware and developing faster models and implementations.
It is worth noting that NPField is a learning-based method and therefore it has no theoretical safety guarantees. Therefore, the obtained solution has to be checked for collisions before its execution.
@@
A significant assumption for the MPC problem is that for each dynamic obstacle, we know either a prediction of its future trajectory (NPField-StaticMLP) or a constant direction of its movement. We compensate for the inaccuracy of this assumption via replanning the trajectory on each step. Meaningful directions of future research include integrating neural potential forecasting with advanced motion prediction techniques, uncertainty-aware obstacle models, and joint multi-obstacle architectures that capture interaction effects beyond additive superposition.
Acknowledgments
The study was supported by the Ministry of Economic Development of the Russian Federation (agreement No. 139-15-2025-013, dated June 20, 2025, IGK  [^00] C313925P4B0002).